OpenStudy (anonymous):

Find the exact value of the expression please! sin(tan^-1 2)

OpenStudy (anonymous):



OpenStudy (jdoe0001):

\(\bf tan^{-1}(2)\implies tan^{-1}\left(\cfrac{2}{1}\right)\implies \cfrac{\textit{opposite side}}{\textit{adjacent side}}\implies \cfrac{b}{a}\\ \quad \\ \textit{pythagorean theorem }c^2=a^2+b^2\implies c=\sqrt{a^2+b^2}\\ \quad \\ sin(\theta)=\cfrac{b}{c}\implies \cfrac{\textit{opposite side}}{\textit{hypotenuse}}\)